# Online Resource Allocation via Static Bundle Pricing

### v1

@@ 3.2 The Unconstrained Algorithm
Since UAlg does not take into account item capacities, it may oversell some items. Define by $L_{M,I}(e)$ the number of copies that $\textnormal{{UAlg}}_{M,I}$ sells for item $e$. In the following lemma, proved in Appendix C, we show that UAlg respects the scaled-down capacities in expectation.
@@
For every item $e\in\mathcal{M}$ and any scaling factor $\gamma>1$, there exist $\{0,1\}$ negatively associated random variables $Z_{1},\ldots,Z_{k}$ such that
$L_{M,I}(e)\leq\sum_{i=1}^{k}Z_{i}=Z_{M,I}(e)\quad\text{and}\quad\mathbb{E}_{M,I}\left[L_{M,I}(e)\right]\leq\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\leq\frac{c(e)}{\gamma}.$
4 d-Single-Minded Setting
In this section, we apply our framework to the case in which every buyer demands a single bundle $S_{b}\subseteq\mathcal{M}$ of size at most $d$. We proceed to prove Theorem˜1.1.
Approach.
We follow the general algorithmic framework from Section˜3. We begin by solving the scaled down LP from (Step 1), choosing $\gamma\;=\;e(10d)^{1/B}$. We then construct the randomized menu as in (Step 2).
Analysis.
Lemma 3.4 together with our choice of $\gamma=e(10d)^{1/B}$ immediately yields the following corollary.
@@
For every item $e\in\mathcal{M}$,
$\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\;\leq\;\frac{c(e)}{\gamma}\;=\;\frac{c(e)}{e(10d)^{1/B}}.$
We now show that, for every item $e\in\mathcal{M}$, the probability that its load exceeds its capacity is at most $1/(10d)$. Recall from Corollary 4.1 that
$\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\;\leq\;\frac{c(e)}{\gamma},\qquad\text{where }\gamma=e(10d)^{1/B}.$
Since $Z_{M,I}(e)$ is a sum of negatively associated binary random variables, as shown in the proof of Lemma˜3.4, we apply the standard Chernoff bound. Writing $\mu=\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]$, for every $e\in\mathcal{M}$ we obtain
$\mathbb{P}\!\left[Z_{M,I}(e)\geq c(e)\right]\;\leq\;\left(\frac{e\mu}{c(e)}\right)^{c(e)}\;\leq\;\left(\frac{e}{\gamma}\right)^{c(e)}.$
Substituting our choice of $\gamma$ yields $\mathbb{P}\!\left[Z_{M,I}(e)\geq c(e)\right]\;\leq\;\frac{1}{10d}$ and from union bound we get for every desired set $S$
@@
since $|S|\leq d$. When a buyer arrives, they may prefer some menu entry, but the algorithm may be forced to reject them because one of the items in that bundle has exhausted its capacity. In this case, we say that the buyer is rejected by the algorithm. Let
$\textsc{Blocked}_{M,I}=\{b\in\mathcal{N}:b\in\textnormal{{UAlg}}_{M,I}\text{ and }b\notin\textnormal{{Alg}}_{M,I}\},$
be the set of rejected buyers and $\textnormal{{Alg}}_{M,I}$ the set and the total value of the buyers that our algorithm gets under realization $I$ and menu $M$. We begin with the basic decomposition (note that UAlg contains Alg):
We now prove the following lemma that bounds the lost value due to rejected buyers and leads to our desired approximation ratio.
@@
We have
$\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right]\;\leq\;0.1\;\textnormal{{FracOpt}}_{\gamma}.$
@@
Fix a desired set $S$ with important value $w=w_{S}$ and denote by $\mathcal{A}=\mathcal{A}_{S}$ the set of buyers that desire $S$. Note that a buyer $b$ belongs to $\textsc{Blocked}_{M,I}$ if it belongs to $\textnormal{{UAlg}}_{M,I}$ and at the same time due to limited item capacities $\textnormal{{Alg}}_{M,I}$ cannot accept it. Denote by $\mathcal{E}_{b}$ the latter event where due to capacity constraints $\textnormal{{Alg}}_{M,I}$ cannot accept buyer $b$. It holds that
@@
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}\cap\textsc{Blocked}_{M,I}}v_{b}\right]=\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}v_{b}\mathbbm{1}{\left\{b\in\textsc{Blocked}_{M,I}\right\}}\right]$

### v2

@@ 6.1 The Unconstrained Algorithm and Its Properties
Since UAlg ignores item capacities, we need to bound how many copies of each item are allocated by UAlg. Let $L_{M,I}(e)$ be the random variable that denotes the number of copies of item $e$ allocated by $\textnormal{{UAlg}}_{M,I}$. We define a collection of random variables that (as we will show) upper bound $L_{M,I}(e)$ and allow us to apply concentration inequalities.
Definition 6.2.
For a bundle $S$, let $k_{S}$ be the (deterministic) number of additional copies of $S$ posted at $S$’s important value $w_{S}$ by Algorithm 1 ($k_{S}=\left\lfloor x_{S,w_{S}}\right\rfloor$, if line 6 is executed, and $k_{S}=0$ otherwise), and let $Y_{S}\in\{0,1\}$ be the (possibly random) indicator that one additional copy of $S$ is posted at price $w_{S}$ (line 8). We define the following random variables for each bundle $S$ and each item $e$:
$Z_{M,I}(S)\coloneqq\sum_{b\in\mathcal{N}_{S}}\mathbbm{1}{\left\{v_{b}>w_{S}\right\}}+k_{S}+Y_{S}\cdot\mathbbm{1}{\left\{\exists b\in\mathcal{N}_{S}\,:\,v_{b}=w_{S}\right\}}\quad\text{and}\quad Z_{M,I}(e)\coloneqq\sum_{S:e\in S}Z_{M,I}(S).$
The following, proven in Appendix C, shows that UAlg respects item capacities in expectation.
@@
For every item $e\in\mathcal{M}$ and any scaling factor $\gamma>1$, it holds that
$L_{M,I}(e)\leq Z_{M,I}(e)\quad\text{and}\quad\mathbb{E}_{M,I}\left[L_{M,I}(e)\right]\leq\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\leq c(e)/\gamma\,.$
In addition, $Z_{M,I}(e)$ is a sum of negatively associated (NA) random variables.
Our analysis upper bounds the lost value from buyers that belong to UAlg but not to Alg. Let
$\textsc{Blocked}_{M,I}=\left\{b\in\mathcal{N}:b\in\textnormal{{UAlg}}_{M,I}\text{ and }b\notin\textnormal{{Alg}}_{M,I}\right\},$ (1)
be the set of not accepted buyers in UAlg and let $\textnormal{{Alg}}_{M,I}$ denote the set and the total value of the buyers that our algorithm accepts under realization $I$ and menu $M$. It holds that
7 d-Single-Minded Combinatorial Auctions: The Proof of Theorem 2.1
Next, we apply our framework to the case where every buyer demands a single bundle $S_{b}\subseteq\mathcal{M}$ of size at most $d$ and prove Theorem 2.1. To this end, we follow the general framework presented in Section 6. We begin by solving the $\gamma$-scaled down EA-LP, with $\gamma=e(10d)^{1/B}$, construct the randomized menu $M$ as in Algorithm 1 and apply the online algorithm presented in Step 3.
Competitive Ratio.
Lemma 6.3 and the choice of $\gamma=e(10d)^{1/B}$ immediately imply that:
@@
For every item $e\in\mathcal{M}$, $\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\leq c(e)/\gamma=c(e)/e(10d)^{1/B}\,.$
We first show that the probability that a bundle $S$ allocated by the unconstrained mechanism is blocked due to item capacities is at most $0.1$. Since $Z_{M,I}(e)$ is a sum of negatively associated binary random variables and upper bounds $L_{M,I}(e)$, as shown in Lemma 6.3, we apply the standard Chernoff bound to $Z_{M,I}(e)$. Using $\mu_{e}=\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\leq c(e)/\gamma$, for every $e\in\mathcal{M}$, we obtain that
$\mathbb{P}\!\left[Z_{M,I}(e)\geq c(e)\right]\;\leq\;\left(\frac{e\mu_{e}}{c(e)}\right)^{c(e)}\;\leq\;\left(\frac{e}{\gamma}\right)^{c(e)}\;\leq\;\frac{1}{10d}\,$,
by our choice of $\gamma$. Applying a union bound and $|S|\leq d$, we get that for every desired bundle $S$,
@@
When a buyer arrives, they may prefer some menu entry (so they are accepted by UAlg), but they may not be accepted by the actual algorithm because one of the items in their bundle has exhausted its capacity. Then, we say that the buyer is blocked by the algorithm (see also (1)). The following bounds the expected social welfare lost due to blocked buyers and implies Theorem 2.1.
@@
We have that $\quad{\displaystyle\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right]\;\leq\;0.1\;\textnormal{{FracOpt}}_{\gamma}\,.}$
@@
The event that a buyer $b$ is blocked depends, in a delicate way, on their value $v_{b}$ and the realized values of all other buyers (the realized values of all buyers comprise the instance $I$, both the buyers’ arrival order determined by the almighty adversary and the algorithm’s set of accepted buyers when $b$ arrives depend on $I$) and the randomized construction of the algorithm’s menu $M$. The proof decouples the above sources of randomness via carefully selected events.
We fix a bundle $S$ with important value $w=w_{S}$. We let $\mathcal{A}=\mathcal{N}_{S}$ be the set of buyers interested in $S$. We let $\mathcal{E}_{b}$ denote the event that buyer $b$ is not accepted due to exhausted item capacities.
We upper bound the expected welfare lost due to blocked buyers in $\mathcal{A}$ by analyzing separately the loss due to blocked buyers with realized values larger than (resp. equal to) the important value $w$, denoted $T_{>w}$ and $T_{=w}$, respectively. Formally, we let
@@
$T_{>w}\coloneqq\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}v_{b}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\right]\text{ and }T_{=w}\coloneqq w\cdot\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}=w\right\}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\right]$
It suffices to show that $T_{>w}+T_{=w}\leq 0.1\,\sum_{v>w}v\,x_{S,v}+0.1\,w\,x_{S,w}$. Then, the lemma follows by applying linearity of expectation over all desired bundles $S$.
For a fixed bundle $S$ and each item $e\in S$, we distinguish between the demand for $e$ due to allocated copies of bundles different from $S$, denoted $Z^{\neg S}_{M,I}(e)\coloneqq\sum_{T\neq S:\ e\in T}Z_{M,I}(T)$, and the demand for $e$ due to allocated copies of $S$, denoted $Z_{M,I}(S)$ (see also Definition 6.2). Then, $Z_{M,I}(e)=Z^{\neg S}_{M,I}(e)+Z_{M,I}(S)$. We observe that $Z^{\neg S}_{M,I}(e)$ depends only on the realized values of buyers not in $\mathcal{A}$ and the randomness in the menu construction for bundles $T\neq S$; it does not depend on $v_{b}$, of any $b\in\mathcal{A}$, or on $Y_{S}$.
Bounding $T_{>w}$.
To apply linearity of expectation in the definition of $T_{>w}$, we fix a buyer $b\in\mathcal{A}$. To decouple $b$’s realized value $v_{b}$ from the events $\mathbbm{1}{\left\{v_{b}>w\right\}}$ and $\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}$ (and these two from each other) in the resulting expectation, we define the following event as a substitute for $\mathcal{E}_{b}$:
$\mathcal{G}_{b}\;\coloneqq\;\left\{\exists e\in S:\ Z^{\neg S}_{M,I}(e)+\!\!\sum_{b^{\prime}\in\mathcal{A}\setminus\{b\}}\!\!\!\mathbbm{1}{\left\{v_{b^{\prime}}>w\right\}}+k_{S}+Y_{S}\cdot\mathbbm{1}{\left\{\exists b^{\prime}\in\mathcal{A}\setminus\{b\}:v_{b^{\prime}}=w\right\}}\,\geq\,c(e)\right\}$
(see Definition 6.2 for the definition of $k_{S}$ and $Y_{S}$). Intuitively, the event $\mathcal{E}_{b}$ (i.e., that $b$ is rejected) implies $\mathcal{G}_{b}$. However, $\mathcal{G}_{b}$ does not depend on the buyers’ arrival order or on $b$’s realized value $v_{b}$ (because it only depends on the realized values of buyers in $\mathcal{A}\setminus\{b\}$ and on the randomness in the menu construction). The following establishes the key property of (and the intuition behind) $\mathcal{G}_{b}$.
Claim 7.3.
$\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\leq\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{G}_{b}\right\}}$.
Proof.
Whenever $v_{b}>w$, buyer $b$ is accepted by $\textnormal{{UAlg}}_{M,I}$ because $M$ includes practically unlimited copies of $S$ at price $w+1$. We show that if additionally $\mathcal{E}_{b}$ occurs, then $\mathcal{G}_{b}$ occurs as well. If $\mathcal{E}_{b}$ occurs, when $b$ arrives there exists some $e\in S$ whose capacity has been exhausted by previously accepted buyers (other than $b$). The contribution of buyers not in $\mathcal{A}$ to the load of $e$ is at most $Z^{\neg S}_{M,I}(e)$. The number of previously accepted buyers from $\mathcal{A}\setminus\{b\}$ is at most $\sum_{b^{\prime}\in\mathcal{A}\setminus\{b\}}\mathbbm{1}{\left\{v_{b^{\prime}}>w\right\}}$ plus at most $k_{S}$ buyers with value $w$ (getting the $k_{S}$ deterministic copies of $S$ with price $w$) plus possibly one additional buyer in $\mathcal{A}\setminus\{b\}$ with value $w$, if $Y_{S}=1$ and such a buyer exists. ∎
Also, by Definition 6.2, for every $e\in S$ and every menu $M$ and instance $I$, we have that
$Z^{\neg S}_{M,I}(e)+\underbrace{\sum_{b^{\prime}\in\mathcal{A}\setminus\{b\}}\!\!\!\mathbbm{1}{\left\{v_{b^{\prime}}>w\right\}}+k_{S}+Y_{S}\cdot\mathbbm{1}{\left\{\exists b^{\prime}\in\mathcal{A}\setminus\{b\}:v_{b^{\prime}}=w\right\}}}_{\leq Z_{M,I}(S)}\,\leq\,Z^{\neg S}_{M,I}(e)+Z_{M,I}(S)=Z_{M,I}(e)$
Therefore, $\mathbb{P}\!\left[\mathcal{G}_{b}\right]\leq\mathbb{P}\!\left[\exists e\in S:\ Z_{M,I}(e)\geq c(e)\right]\leq 0.1$.
Using independence of $\mathcal{G}_{b}$ and $v_{b}$, we obtain that
$\mathbb{E}_{M,I}\left[v_{b}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\right]\leq\mathbb{E}_{M,I}\left[v_{b}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{G}_{b}\right\}}\right]=\mathbb{P}\!\left[\mathcal{G}_{b}\right]\cdot\sum_{v>w}v\,q_{b,v}\leq 0.1\sum_{v>w}v\,q_{b,v}.$
Applying linearity of expectation over all buyers $b\in\mathcal{A}$ and using that $w$ is the important value of $S$ (and thus, all values $v>w$ are tight for $S$), we conclude that
$T_{>w}\ \leq\ 0.1\sum_{b\in\mathcal{A}}\sum_{v>w}v\,q_{b,v}\ =\ 0.1\sum_{v>w}v\,q_{S,v}\ =\ 0.1\sum_{v>w}v\,x_{S,v},$
Bounding $T_{=w}$.
We need to show that the expectation in the definition of $T_{=w}$ is at most $0.1\,x_{S,w}$. We let $p_{S}\coloneqq\mathbb{P}\!\left[Y_{S}=1\right]$ denote the probability that an additional copy of $S$ is posted at price $w$ (in line 8 of Algorithm 1, see also Definition 6.2 for $Y_{S}$). We observe that conditional on $k_{S}>0$, $p_{S}=0$. Hence, we get that $k_{S}+p_{S}\cdot\min\{1,q_{S,w}\}\leq x_{S,w}$, which holds unconditionally.
We let $N_{w}\coloneqq\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}=w\right\}}$ be the random variable that denotes the number of buyers $b\in\mathcal{A}$ with $v_{b}=w$. A difficulty in bounding $T_{=w}$ arises from the need to consider all buyers $b\in\mathcal{A}$ with $v_{b}=w$ together, so that we can apply the upper bound of $k_{S}+Y_{S}\cdot\mathbbm{1}{\left\{N_{w}\geq 1\right\}}$ on the number of such buyers accepted by UAlg (see also Definition 6.2 for $k_{S}$). Hence, we define the following event as a substitute of the event that some buyers in $b\in\mathcal{A}$ with $v_{b}=w$ are not accepted:
$\mathcal{F}_{S}\coloneqq\left\{\exists e\in S:\ Z^{\neg S}_{M,I}(e)+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}+k_{S}\ \geq\ c(e)\right\}$
We observe that $\mathcal{F}_{S}$ does not depend on the buyers’ arrival order and that its occurrence is negatively associated with $N_{w}$ (because each buyer $b\in\mathcal{A}$ has either $v_{b}\leq w$ or $v_{b}>w$, but not both). Whenever there exists a buyer $b\in\mathcal{A}$ with $v_{b}=w$ not accepted by the algorithm (i.e., the corresponding event $\mathcal{E}_{b}$ occurs), the event $\mathcal{F}_{S}$ occurs as well. Moreover, for every item $e\in S$, every realized instance $I$ and every realized menu $M$, the sum accounting for the load of $e$ in the definition of $\mathcal{F}_{S}$ does not exceed $Z_{M,I}(e)$. Therefore, $\mathbb{P}\!\left[\mathcal{F}_{S}\right]\leq\mathbb{P}\!\left[\exists e\in S:\ Z_{M,I}(e)\geq c(e)\right]\leq 0.1$.
The total number of buyers $b\in\mathcal{A}$ with $v_{b}=w$ accepted by UAlg (such buyers $b$ may trigger an event $\mathcal{E}_{b}$ in the definition of $T_{=w}$) is at most $k_{S}$ (via the $k_{S}$ deterministic copies of $S$ posted at price $w$ in Algorithm 1, line 6) plus at most one additional such buyer, in case where $Y_{S}=1$ and $N_{w}\geq 1$. Hence, for every realized instance $I$ and every realized menu $M$, we have that:
$\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}=w\right\}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\ \leq\ \Big(k_{S}+Y_{S}\cdot\mathbbm{1}{\left\{N_{w}\geq 1\right\}}\Big)\cdot\mathbbm{1}{\left\{\mathcal{F}_{S}\right\}}.$
Taking expectations and using that the realization of $Y_{S}$ depends only on the randomness of an independent coin flip (for bundle $S$) in Algorithm 1, line 8 (so $Y_{S}$ does not depend on the realized instance $I$ or on $\mathcal{F}_{S}$) and that $\mathbbm{1}{\left\{\mathcal{F}_{S}\right\}}$ and $\mathbbm{1}{\left\{N_{w}\geq 1\right\}}$ are negatively associated,) we obtain that:
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}=w\right\}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\right]\ \leq\ \Big(k_{S}+p_{S}\cdot\mathbb{E}_{M,I}\left[\mathbbm{1}{\left\{N_{w}\geq 1\right\}}\right]\Big)\cdot\mathbb{P}\!\left[\mathcal{F}_{S}\right]$ (2)
We observe that $\mathbb{E}_{M,I}\left[\mathbbm{1}{\left\{N_{w}\geq 1\right\}}\right]\leq\min\{1,q_{S,w}\}$, since $\mathbbm{1}{\left\{N_{w}\geq 1\right\}}\leq 1$ and $\mathbbm{1}{\left\{N_{w}\geq 1\right\}}\leq N_{w}$, with $\mathbb{E}_{M,I}\left[N_{w}\right]=\sum_{b\in\mathcal{A}}\mathbb{P}\!\left[v_{b}=w\right]=q_{S,w}$. Therefore, the right-hand-side of (2) can be upper bounded by $(k_{S}+p_{S}\cdot\min\{1,q_{S,w}\})\cdot\mathbb{P}\!\left[\mathcal{F}_{S}\right]$.
Using that $\mathbb{P}\!\left[\mathcal{F}_{S}\right]\leq 0.1$ and that $k_{S}+p_{S}\cdot\min\{1,q_{S,w}\}\leq x_{S,w}$, we conclude $T_{=w}\leq 0.1\,w\,x_{S,w}$ . ∎
Applying Lemma 7.2, ($\star$6.1) and Lemma 6.1, we conclude the proof of Theorem 2.1:
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;\geq\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\;-\;0.1\,\textnormal{{FracOpt}}_{\gamma}\geq\frac{\textnormal{{FracOpt}}_{\gamma}}{40}\;\geq\;\frac{\textnormal{{FracOpt}}}{40e(10d)^{1/B}}$
8 General Single-Minded Combinatorial Auctions: The Proof of Theorem 2.2
Next, we apply our general algorithmic framework to the case where the buyers are single-minded over arbitrary bundles and prove Theorem 2.2.
The main challenge in the proof of Theorem 2.2 is that in certain (very low probability) instances, the realized value of some buyer may be extremely large, much larger than our fractional benchmark $\textnormal{{FracOpt}}_{\gamma}$. Then, the almighty adversary can arrange that the buyers with extremely large realized value arrive very last, so that it maximizes the probability that they are rejected by the algorithm (while they are accepted by UAlg). To deal with this case, we introduce an auxiliary “small market” mechanism $\textnormal{{Alg}}_{2}$, which posts a single bundle consisting of entire $\mathcal{M}$ at price $2\,\textnormal{{FracOpt}}_{\gamma}$. Our final mechanism randomizes between the auxiliary mechanism $\textnormal{{Alg}}_{2}$ (with probability $2/3$) and our standard “large market” mechanism $\textnormal{{Alg}}_{1}$ obtained from the general framework in Section 6 (with probability $1/3$). For $\textnormal{{Alg}}_{1}$, we solve the $\gamma$-scaled down EA-LP, with $\gamma=e(20m)^{1/(B+1)}$, construct the randomized menu $M$ as in Algorithm 1 and apply the online algorithm presented in Step 3.
As in Section 7, we need to upper bound the expected social welfare lost due to blocked buyers. To this end, we account for the contribution of buyers with extremely large and moderate realized value separately. Formally, we decompose (and upper bound) $\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right]$ as follows:
$\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\,\mathbbm{1}{\left\{v_{b}<2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]+\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]\\ \leq 2\,\textnormal{{FracOpt}}_{\gamma}\cdot\mathbb{E}_{M,I}\left[|\textsc{Blocked}_{M,I}|\right]+\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{N}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$ (3)
Bounding the Number of Blocked Buyers.
We begin by analyzing the first term. The following, whose detailed proof is deferred to Appendix F, bounds the expected number of blocked buyers.
Lemma 8.1.
We have that $\mathbb{E}_{M,I}\left[|\textsc{Blocked}_{M,I}|\right]\;\leq\;1/20$ .
Proof Sketch.
A buyer is blocked only if some item is oversold, hence
$|\textsc{Blocked}_{M,I}|\leq\sum_{e\in\mathcal{M}}(L_{M,I}(e)-c(e))_{+}=\sum_{e\in\mathcal{M}}\sum_{j\geq 1}\mathbbm{1}{\left\{L_{M,I}(e)\geq c(e)+j\right\}}.$
Taking expectations and using Lemma 6.3, a Chernoff bound for negatively associated (NA) variables with $\mu_{e}=\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\leq c(e)/\gamma$, gives that for every item $e\in\mathcal{M}$,
$\mathbb{P}\!\left[L_{M,I}(e)\geq c(e)+j\right]\leq\mathbb{P}\!\left[Z_{M,I}(e)\geq c(e)+j\right]\leq\frac{1}{20m}e^{-j}\,.$
Summing over all $j\geq 1$ yields $\mathbb{E}_{M,I}\left[(L_{M,I}(e)-c(e))_{+}\right]\leq 1/(20m)$. Summing over all items $e\in\mathcal{M}$ concludes the proof of the lemma. ∎
Plugging Lemma 8.1 into (3) and using ($\star$6.1) and Lemma 6.1, we obtain that:
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;\geq\;\frac{\textnormal{{FracOpt}}_{\gamma}}{40}-\sum_{b\in\mathcal{N}}\mathbb{E}_{M,I}\left[v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]\,.$ (4)
Dealing with Extremely Large Values.
The negative term of (4) corresponds to buyers with extremely large realized values. To deal with them, we introduce auxiliary mechanism $\textnormal{{Alg}}_{2}$, which posts a single menu entry consisting of all items, priced at $2\,\textnormal{{FracOpt}}_{\gamma}$. We distinguish two cases.
Case 1: $\mathbb{P}\!\left[\exists b:v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right]>1/2$. Then, simply posting one copy of entire $\mathcal{M}$ at price $2\,\textnormal{{FracOpt}}_{\gamma}$ yields an expected social welfare larger than $\textnormal{{FracOpt}}_{\gamma}$, ensuring an $O(\gamma)$-approximation directly.
Case 2: $\mathbb{P}\!\left[\exists b:v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right]\leq 1/2$. The following shows that in this case, the expected social welfare of $\textnormal{{Alg}}_{2}$ is at least half of the negative term of (4).
Lemma 8.2.
If  $\mathbb{P}\!\left[\exists b:v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right]\leq 1/2$, then  $\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{2}\right]\geq\frac{1}{2}\,\sum_{b\in\mathcal{N}}\mathbb{E}_{M,I}\left[v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right\}}\right].$
Proof.
For every fixed realized instance $I$, we have that: